OpenStudy (anonymous):

Need help understanding this method and how to slove :/ Solve each of the following problems using a system of equations. Show all of your work and state your solution in a complete sentence. Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each? (2 points) Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (2 points) Follow the 5 steps below to complete this problem. (4 points) Step 1: Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following: Drive a boat Drive a jet ski Step 2: Select a current speed of the water in mph. Step 3: Select the number of hours (be reasonable please) that your friend or family member drove the boat or jets ski against the current speed you chose in step 2. Step 4: Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling with the current). Step 5: Write out the word problem you created and calculate how fast your friend or family member was traveling in still water. Round your answer to the nearest mph.

OpenStudy (anonymous):

Tony and Linda have a combines age of 56 so....t + b = 56 Belinda is 8 more then twice Tony's age so....b = 2t + 8. Now we have our two equations. We know that b = 2t + 8, so sub 2t + 8 in for b in the 1st equation. t + b = 56 t + 2t + 8 = 56 combine like terms 3t + 8 = 56 subtract 8 from each side to get 3t by itself 3t = 56 - 8 3t = 38 divide by 3 to get t by itself t = 48/3 t = 16.....so Tony's age is 16 now sub 16 in for t in the 1st equation t + b = 56 16 + b = 56 can you finish and solve for Belinda's age ?

OpenStudy (anonymous):

oops...mistake in there...should be 3t = 48

OpenStudy (anonymous):

Oh i see owo

OpenStudy (anonymous):

b= 40?

OpenStudy (anonymous):

the second problem can be solved by substitution as well.... a + s = 150 45a + 28.5s = 4770 can you solve this one ?

OpenStudy (anonymous):

b = 40 in your last question.....and t = 16

OpenStudy (anonymous):

do you know how to do the second problem ?